OpenStudy (anonymous):

Find the area bounded by x+xy^2-1=0 and y^2 -2x=0.

OpenStudy (amistre64):

looks to me like you might want to define 2 function of y, then integrate between them

OpenStudy (amistre64):

x+xy^2-1=0 x(1+y^2) = 1 x = 1/(1+y^2) y^2 -2x=0. y^2 = 2x (y^2)/2 = x

OpenStudy (anonymous):

Thank you for the reply. And I saw that far but the problem was getting correct parameters and then correctly integrating the equations.My teacher never went too far into that

OpenStudy (amistre64):

can you calculate the intersections of these? itll give us an interval to work across

OpenStudy (amistre64):

 the basic process will be \[\int_{a}^{b}~f-g\]